OpenStudy (anonymous):

Simplify: Square root x^3 divided by Cube root x^4

OpenStudy (anonymous):

how else could you write this out?

OpenStudy (anonymous):

http://imgur.com/ckiMgdn this is a picture of the question

OpenStudy (anonymous):

Yes, I know how to write it out. Do you know how else to write it out?

OpenStudy (mertsj):



OpenStudy (anonymous):

just like that^ cheetos

OpenStudy (anonymous):

Then it's a matter of subtracting the second exponent from the first.

OpenStudy (anonymous):

yes wrote it like that tan you subtract the exponets right?

OpenStudy (anonymous):

uh huh!

OpenStudy (anonymous):

so it becomes x^-1/-1 ?

OpenStudy (anonymous):

how did you get that? did you subtract 4/3 from 3/2?

OpenStudy (anonymous):

which turns into x^1 because negative by negative becomes positive right?

OpenStudy (anonymous):

3/2 - 4/3 = ?

OpenStudy (mertsj):

\[\frac{3}{2}-\frac{4}{3}=\frac{9}{6}-\frac{8}{6}=\frac{1}{6}\]

OpenStudy (anonymous):

^, this guy isn't helping you, he's doing your work for you!

OpenStudy (anonymous):

? i thought 3-4 is negative 1 because 4 is bigger

OpenStudy (anonymous):

just copy verbatim what Mertsj is writing and you're on your way!

OpenStudy (anonymous):

oh you have to have same denominator

OpenStudy (anonymous):

you need to have a common denominator when subtracting fractions.

OpenStudy (mertsj):

You can't subtract fractions unless the denominators are the same.

OpenStudy (anonymous):

Ok thanks sean and mertsj

OpenStudy (anonymous):

don't mention it.

OpenStudy (mertsj):

yw